4. Find the currents 11, 12, and 13 and the terminal voltage of the ε2 battery -0.3 Ω r2-0.2 Ω et 0.96 52 3
Answer #1

at point a


applying kirchoff junction law

I1 + I2 = I3.......(1)


for upper loop


from kirchoff voltage law


E1 - I1*r1 - E2 + I2*r2 = 0


6 - I1*0.3 - 5 + I2*0.2 = 0.........(2)


for lower loop

E2 - I2*r2 - I3*R = 0

5 - I2*0.2 - I3*0.96 = 0 .......(3)


solving 1 , 2, 3

I1 = 4 A


I2 = 1 A


I3 = 5 A